OpenStudy (anonymous):

Please Help I suck at algebra!!!!!!! Which of the following lists all x-intercepts of the graph of the equation y = x3 – 4x? (1 point) a.) (0, 0), (4, 0) b.) (–2, 0), (0, 0), (2, 0) c.) (–2, 0), (2, 0) d.) (–4, 0), (0, 0), (4, 0)

OpenStudy (anonymous):

Is that 3 an exponent?

OpenStudy (anonymous):

yes sorry about that.

OpenStudy (anonymous):

x intercepts happen when y = 0. First you factor the equation into y = x(x^2 - 4) Now set that = 0 x(x^2 - 4) = 0 when x = 0 or when x^2 - 4 = 0 x^2 - 4 = 0 when x^2 = 4 therefore when x= 2 or x = -2 So the x-intercepts occur at (0,0), (2,0) and (-2,0)

OpenStudy (anonymous):

Okay thank you so much! :)